# Xiang Fan1,2    Sharath Girish2    Vivek Ramanujan1    Chaoyang Wang2    Ashkan Mirzaei2    Petr Sushko1    Aliaksandr Siarohin2    Sergey Tulyakov2    Ranjay Krishna1

### v1

@@ 1 Introduction
When trained on raw internet-scale video, video diffusion models internalize strong 3D priors [snapvideo, sv3d, li2024director3d]. They can synthesize long, coherent camera motions and maintain scene layout without any explicit geometry [genie3]. Yet, by default, they offer little explicit 3D control. To deploy them in applications such as virtual/augmented reality, film production, robotics, or autonomous driving, they need to be controllable. Many applications would become possible if we could control how the generator’s camera moves in space and in time. Naturally, this motivation has driven multiple sub-fields of computer vision to develop specialized solutions. Camera redirection now exists for multi-view static [gao2024cat3d, zhou2025stable] and dynamic Novel View Synthesis (NVS) [syncammaster, wang20254real], text-to-video (T2V) [cameractrl, dav24] and image-to-video (I2V) [camco] with camera control, and even video-to-video (V2V) [bai2025recammastercameracontrolledgenerativerendering].
However, existing approaches are fragmented along task, architecture, and data. Methods designed for multi-view static NVS focus on reconstructing a single scene from sparse views [zhou2025stable], achieving strong 3D consistency but only at a fixed timestamp, and thus cannot handle dynamic videos. Camera-control T2V/I2V models convert text or a single image plus a camera trajectory and generate a moving video, but their architectures cannot ingest full input videos [cameractrl]. V2V redirection models accept a source video and re-render it from a new camera path at matched timestamps, yet they typically cannot operate in the multi-view setting [bai2025recammastercameracontrolledgenerativerendering]. Some works rely on explicit geometric representations (depth maps, point clouds, or other 3D/4D fields) for consistency [yu2025trajectorycrafter, ren2025gen3c, yu2024viewcrafter, wu2024reconfusion], rather than exploiting the implicit 3D priors already present in video models.
@@
To realize this unified 4D formulation, we adopt a Diffusion Transformer (DiT) [dit] backbone that naturally handles a variable number of input tokens. We tokenize each frame into a set of video tokens, concatenate tokens from all available inputs (images, views, or frames), and condition generation on this sequence. DiTs already support temporal reasoning via spatio-temporal Rotary Positional Embeddings (3D RoPE), which encode $(x,y,t)$ for each token. Prior works typically injects camera information by either encoding poses with a pose encoder or by mapping them to Plücker ray embeddings and then applying 3D RoPE [su2024roformer] to both video and camera-conditioned tokens. This entangles camera pose $\mathbf{p}\mathbf{p}$ and time $t$ in a single positional embedding space, making it difficult for the model to learn 3D structure independently of temporal dynamics and often leading to overfitting to seen trajectories.
In contrast, OmniView explicitly disentangles space and time. We represent each token’s camera only to as ase Plücker nd apply spatial 2D RoPE only to these Plucker features, then concatenate the them channel-wise with the corresponding video token. Time is encoded separately via temporal RoPE on the video tokens. This design cleanly separates camera geometry from temporal evolution, while still allowing the DiT to jointly attend over all tokens. Combined with the corresponding video token. Time is encoded separately via temporal RoPE on the video tokens. This design cleanly separates camera geometry from temporal evolution, while still allowing the DiT to jointly attend over all tokens. Combined with the variable-token design, this lets OmniView flexibly ingest arbitrary combinations of frames, views, and timestamps and thereby support a wide range of 4D inputs, under a single architecture. We thenthe variable-token design, this lets OmniView flexibly ingest arbitrary combinations of frames, views, and timestamps and thereby support a wide range of 4D inputs under a single architecture. We then devise a joint training strategy that mixes heterogeneous 3D datasets, each correspondconditioned on text and images. OmniView, 3D-consistently matches or outperforms specialized baselines, producing to different task configurations (multi-view static, dynamic, T2V/I2V with camera control, V2V redirection), so that the model learns shared geometric priors across them.
We extensively evaluate OmniView on static and dynamic NVS benchmarks with monocular and multi-view inputs, as well as camera-control tasks conditioned on text and images. OmniView consistently matches or outperforms specialized baselines, producing high-fidelity, 3D-consistent videosand camera-conditioned tokens. With up to 33% increase in SSIM scores in multiview NVS LLFF dataset, 60% in dynamic NVS Neural 3D Video benchmark, 20% in static camera control on RE-10K, 4x reduction in camera errors in text-conditioned video generation, OmniView demonstrates strong 3D consistency and fidelity across tasks, and generalization to input configurations not seen during training. Ablations validate the benefit of our RoPE design.
@@ 2 Related Work
The emergence of powerful text-to-video diffusion models [blattmann2023stable, videocrafter2, cogvideox, snapvideo, hunyuanvideo] has fueled extensive research on conditioning generated videos with additional controls, such as camera parameters [direct_a_video, vd3d, cami2v, zheng2025vidcraft3, liu2023zero1to3, shi2023zero123plus, vanhoorick2024gcd, animatediff, MotionCtrl, watson2024controlling]. Early camera-control approaches integrate extrinsic camera information as part of the diffusion model’s conditioning, either through tailored encoders or numeric input channels. Models like MotionCtrl [MotionCtrl] and CameraCtrl [cameractrl] encode camera poses or trajectories to enable user-directed viewpoint changes throughout video generation, but often require specific paired training data and show limited generalization when camera motions deviate from training regimes. Other strategies bypass model retraining by employing 3D geometric cues, for example warping frames using estimated depth to match new camera placements and feeding these as priors during the denoising process [NVS_Solver, camtrol], though these methods face a trade-off between enforcing geometric consistency and visual fidelity.
@@
Generating unseen viewpoints from posed images or videos has advanced significantly in recent years [nerfstudio, mildenhall2021nerf, kerbl3Dgaussians, barron2022mipnerf360, yang2023emernerf, Yu2024GOF, Li_2023_CVPR, adaptiveshells2023, Huang2DGS2024, barron2021mipnerf, mueller2022instant, luiten2023dynamic, duan:2024:4drotorgs]. Conventional novel view synthesis (NVS) frameworks leverage volumetric or Gaussian-based scene representations; meanwhile, feed-forward architectures [yu2020pixelnerf, mvsnerf, zhou2023nerflix, tang2024lgm, jin2024lvsmlargeviewsynthesis, wang2021ibrnet, pixelsplat, chen2025mvsplat, ren2024scube] aim to directly predict target views from sparse or multi-view input, but usually struggle in generalization tasks or under challenging domain shifts. Some recent works attempt to harness image/video generative models to infuse prior knowledge and regularize deficits of view synthesis, as in ReconFusion [wu2024reconfusion] and CAT3D [gao2024cat3d]. However, these strategies tend to be slow due to per-scene optimization and often depend on robust inter-view alignment, as seen in ReconX [liu2024reconx] and ViewCrafter [yu2024viewcrafter], which become error-prone in the presence of thin or ambiguous structures. Relatedly, the video-to-video generation field [wang2018video, wang2019few, mallya2020world, rerender, videop2p, wang2024your, chen2024follow, zhou2024upscale, xu2024videogigagan] explores producing temporally consistent and controllable video outputs under various manipulation and conditioning tasks. Techniques such as GCD [vanhoorick2024gcd], Recapture [recapture], GS-DiT [gs-dit], DAS [das], and recent 4D-consistent pipelines [ren2025gen3c, yu2025trajectorycrafter, trajattn] exploit geometric and dynamic scene information, such as tracked 3D points [cotracker, spatialtracker], to condition or align the generative models, either via simulation or real-world sequences. These approaches enable synchronizing generated outputs across multiple cameras or time but are typically constrained by how accurately dynamic scene content can be retrieved or tracked. Some works tackle 4D and multi-view video generation by training generative models directly on synchronized video collections [sv4d, wang20254real, wang20254realv2, syncammaster, cat4d, li2024vivid], or by reconstructing an explicit scene representation first and then rendering views [monst3r, megasam, liu2024reconx, sun2024dimensionx].
@@
Ensuring frame-to-frame and cross-view temporal or geometric consistency remains a critical challenge in video synthesis. Early efforts used 3D point clouds or height maps derived from input or generated content to guide learning and enforce consistency [mallya2020world, deng2024streetscapes]. Others [kuang2024collaborative] propagate consistency across parallel generated video streams but may lose coherence when scene elements leave the views of all sequences. Latent feature histories have also been used to improve consistency for streaming or autoregressive video generation approaches [streamingt2v], though explicit, interpretable 3D control remains an open research direction.
@@ 3.1 Preliminary: Video Diffusion Models
Our framework builds on the popular architecture used in state-of-the-art text-to-video diffusion models [cogvideox, cosmos, hunyuanvideo, wan2025wan], which combine a 3D Variational Auto-Encoder (VAE) with a Diffusion Transformer (DiT) [dit]. The VAE spatially and temporally compresses input videos into low-resolution latents that serve as compact representations for diffusion modeling. The diffusion process follows the rectified flow formulation [liu2022flow], where the model learns to transform noise into coherent video latents through velocity prediction.
@@ 3.2 Network architecture
A naive strategy for injecting camera tokens $\mathbf{c}_{xyt}$ is to simply concatenate or add them to the corresponding video tokens $\mathbf{z}_{xyt}$. A similar approach is adopted in [bai2025recammastercameracontrolledgenerativerendering], where a camera encoder produces a 12-dimensional pose embedding that is added to the video tokens in every DiT block. However, this formulation entangles the spatial location of the camera and the temporal position of the corresponding frame within the video, as discussed below.
Disentangle camera and temporal position embeddings. Video DiTs use 3D Rotary Positional Embeddings (RoPE) [su2024roformer] to encode the spatial-temporal positions $(x,y,t)$ of video tokens. Specifically, RoPE applies a frequency-based rotation on the keys and values of a token:
@@ 3.3 Training setup
To enable the model to accommodate different input configurations across a variety of tasks, we leverage a diverse collection of existing 3D/4D datasets encompassing multiple data types. Stereo4D [jin2025stereo4d] comprises stereo videos with camera-pose annotations; however, because per-video stereo baselines are unavailable, we use a single view with its corresponding pose, effectively treating it as a monocular video dataset. We include RE10K [zhou2018re10k] and DL3DV [ling2024dl3dv] as multi-view image NVS datasets. For image/video NVS, we additionally use the synthetic Syncammaster [syncammaster] and Recammaster [bai2025recammastercameracontrolledgenerativerendering] datasets, featuring multiple time-synchronized static cameras and dynamic camera trajectories, respectively. An overview of the datasets and the tasks each dataset supports is provided in Tab. 1. We primarily target context and target configurations with $1$, 33, 55, or $10$ latent frames, and up to $3$ views for the context input. Owing to the variation in input views and frames across training iterations, we demonstrate in Fig. 5 that our model generalizes to longer video sequences and a greater number of views. Moreover, despite not training directly on the multi-view video NVS task (e.g., $3{\times}3\rightarrow 1{\times}3$), the model generalizes to it naturally, as it can be viewed as a composition of the multi-view image NVS and monocular video NVS tasks.
@@ 4.1 Experimental Setup
Implementation Details. As discussed in Sec. 3.3, we train OmniView on a mixture of the datasets provided in Tab. 1. We choose the Wan 2.1 1.1B model [wan2025wan] as the base architecture for all our experiments unless mentioned otherwise. Each iteration takes in a set of source views, target views and their corresponding cameras as well as their timestamps in the real world. We train the model for 40K iterations over 32 H100 GPUs. We linearly warmup the learning rate for 3K iterations upto $0.0001$ with a batch size of 64. During the warmup stage, we exclusively train on the multiview static task to quickly update the parameters corresponding to the Plucker ray maps while also allowing the model to quickly adapt to camera conditioning.
@@
TrajectoryCrafter [yu2025trajectorycrafterredirectingcameratrajectory] 24.77 3.59 21.36 3.60 14.37 4.00 24.92 2.64 23.11 2.05 21.70 3.18
ReCamMaster [bai2025recammastercameracontrolledgenerativerendering] 10.18 2.60 11.47 2.19 8.85 3.39 16.49 2.82 13.49 2.37 12.09 2.67
Gen3C [ren2025gen3c] 28.38 1.88 42.55 2.04 129.17 2.62 21.80 1.39 25.72 1.71 49.52 1.93
@@
Evaluation and Baselines. We comprehensively evaluate OmniView on a wide variety of 3D/4D tasks. We broadly group multiple tasks under: a) Monocular Video NVS (Tab. 2) b) Multi-view Image and Video NVS (Tab. 3) and c) T2V/I2V + Camera Control. We mainly perform comparisons with closely related generative view-synthesis works of RecamMaster [bai2025recammastercameracontrolledgenerativerendering], TrajectoryCrafter [yu2025trajectorycrafter], and GEN3C [ren2025gen3c] on the Monocular Video NVS as well as the I2V + Camera Control task. For Multi-view Image NVS, GEN3C and SEVA [zhou2025stable] are our primary baselines which can take multiple view inputs. We visualize the results from our model in Fig. 3 and comparisons in the supplementary. We use the reconstruction metrics of PSNR, SSIM, LPIPS for datasets with GT views available. We additionally evaluate the quality of the camera trajectory via Rotation (RotErr) and Translation Error (TrErr) as proposed by [cameractrl]. We use MegaSaM [megasam] for estimating camera trajectories from generated videos when comparing with the GT trajectories. We additionally compute the metric scale for each scene similar to [ac3d] and scale the camera translation vectors accordingly. We find this necessary for monocular view cases which have inherent scale ambiguity.
@@ 4.2 Monocular Video NVS
For this task, we extract camera trajectories for 45 real-world videos from DAVIS [Huang_2016_davis]. For each video, we use 4 predefined trajectory paths from RecamMaster evaluations and 1 new trajectory which is a spiral path. Results are summarized in Tab. 2. Averaged over all trajectories, we match the performance of SOTA approaches of [bai2025recammastercameracontrolledgenerativerendering, yu2025trajectorycrafter, ren2025gen3c] in terms of camera error metrics despite being trained on a large diversity of tasks. We additionally visualize qualitative results in the supplementary. We see that we obtain high fidelity generations compared to other works with fewer motion artifacts. We maintain high temporal synchronization compared to the input condition highlighting the capability of the model to accurately extract appearance information from conditioning views while disentangling camera and time conditions.
Beyond camera trajectory error metrics, we also compare with [yu2025trajectorycrafter, ren2025gen3c] on the Neural 3D Video (N3DV) dataset [li2022neural] which consists of multi-view static cameras capturing a dynamic scene. We choose the first view as the test and choose one training view with significant disparity from the test. Compared to prior works, we obtain high quality reconstructions which are well aligned with the GT achieving better PSNR, SSIM and LPIPS. Note that our conditioning is largely implicit via encoded latents and we do not use any form of explicit 3D supervision such as depth or point clouds apart from the metric scene scale. Despite being trained on 122K camera trajectories, RecamMaster fails for this setting of static target camera and tends to implicitly induce camera motion. This highlights the generalizability issue of using 3D spatiotemporal RoPE on camera conditions as discussed in Sec. 3 where the model fails to disentangle camera and time resulting in poor out-of-distribution trajectory performance.

### v2

@@ 1 Introduction
When trained on raw internet-scale video, video diffusion models internalize strong 3D priors [44, 58, 34]. They can synthesize long, coherent camera motions and maintain scene layout without any explicit geometry [48]. Yet, by default, they offer little explicit 3D control. To deploy them in applications such as virtual/augmented reality, film production, robotics, or autonomous driving, they need to be controllable. Many applications would become possible if we could control how the generator’s camera moves in space and in time. Naturally, this motivation has driven multiple sub-fields of computer vision to develop specialized solutions. Camera redirection now exists for multi-view static [18, 91] and dynamic Novel View Synthesis (NVS) [5, 61], text-to-video (T2V) [21, 78] and image-to-video (I2V) [74] with camera control, and even video-to-video (V2V) [4].
However, existing approaches are fragmented along task, architecture, and data. Methods designed for multi-view static NVS focus on reconstructing a single scene from sparse views [91], achieving strong 3D consistency but only at a fixed timestamp, and thus cannot handle dynamic videos. Camera-control T2V/I2V models convert text or a single image plus a camera trajectory and generate a moving video, but their architectures cannot ingest full input videos [21]. V2V redirection models accept a source video and re-render it from a new camera path at matched timestamps, yet they typically cannot operate in the multi-view setting [4]. Some works rely on explicit geometric representations (depth maps, point clouds, or other 3D/4D fields) for consistency [83, 51, 85, 70], rather than exploiting the implicit 3D priors already present in video models.
@@
To realize this unified 4D formulation, we adopt a Diffusion Transformer (DiT) [49] backbone that naturally handles a variable number of input tokens. We tokenize each frame into a set of video tokens, concatenate tokens from all available inputs (images, views, or frames), and condition generation on this sequence. DiTs already support temporal reasoning via spatio-temporal Rotary Positional Embeddings (3D RoPE), which encode $(x,y,t)$ for each token. Prior works typically injects camera information by either encoding poses with a pose encoder or by mapping them to Plücker ray embeddings and then applying 3D RoPE [53] to both video and camera-conditioned tokens. This entangles camera pose $\mathbf{p}\mathbf{p}$ and time $t$ in a single positional embedding space, making it difficult for the model to learn 3D structure independently of temporal dynamics and often leading to overfitting to seen trajectories.
In contrast, OmniView explicitly disentangles space and time. We represent each token’s camera pose as Plücker rays and apply spatial 2D RoPE only to these Plücker features, then concatenate them channel-wise with the corresponding video token. Time is encoded separately via temporal RoPE on the video tokens. This design cleanly separates camera geometry from temporal evolution, while still allowing the DiT to jointly attend over all tokens. Combined with the variable-token design, this lets OmniView flexibly ingest arbitrary combinations of frames, views, and timestamps and thereby support a wide range of 4D inputs under a single architecture. We then devise a joint training strategy that mixes heterogeneous 3D datasets, each corresponding to different task configurations (multi-view static, dynamic, T2V/I2V with camera control, V2V redirection), so that the model learns shared geometric priors across them.
We extensively evaluate OmniView on static and dynamic NVS benchmarks with monocular and multi-view inputs, as well as camera-control tasks conditioned on text and images. OmniView consistently matches or outperforms specialized baselines, producing high-fidelity, 3D-consistent videos. With up to 33% increase in SSIM scores in multiview NVS LLFF dataset, 60% in dynamic NVS Neural 3D Video benchmark, 20% in static camera control on RE-10K, 4x reduction in camera errors in text-conditioned video generation, OmniView demonstrates strong 3D consistency and fidelity across tasks, and generalization to input configurations not seen during training. Ablations validate the benefit of our RoPE design.
@@ 2 Related Work
The emergence of powerful text-to-video diffusion models [9, 12, 80, 44, 17] has fueled extensive research on conditioning generated videos with additional controls, such as camera parameters [77, 3, 89, 90, 39, 52, 57, 20, 66, 68]. Early camera-control approaches integrate extrinsic camera information as part of the diffusion model’s conditioning, either through tailored encoders or numeric input channels. Models like MotionCtrl [66] and CameraCtrl [21] encode camera poses or trajectories to enable user-directed viewpoint changes throughout video generation, but often require specific paired training data and show limited generalization when camera motions deviate from training regimes. Other strategies bypass model retraining by employing 3D geometric cues, for example warping frames using estimated depth to match new camera placements and feeding these as priors during the denoising process [81, 23], though these methods face a trade-off between enforcing geometric consistency and visual fidelity.
@@
Generating unseen viewpoints from posed images or videos has advanced significantly in recent years [55, 46, 29, 7, 76, 86, 35, 67, 24, 6, 47, 42, 16]. Conventional novel view synthesis (NVS) frameworks leverage volumetric or Gaussian-based scene representations; meanwhile, feed-forward architectures [82, 11, 92, 56, 26, 63, 10, 14, 50] aim to directly predict target views from sparse or multi-view input, but usually struggle in generalization tasks or under challenging domain shifts. Some recent works attempt to harness image/video generative models to infuse prior knowledge and regularize deficits of view synthesis, as in ReconFusion [70] and CAT3D [18]. However, these strategies tend to be slow due to per-scene optimization and often depend on robust inter-view alignment, as seen in ReconX [38] and ViewCrafter [85], which become error-prone in the presence of thin or ambiguous structures. Relatedly, the video-to-video generation field [65, 64, 43, 79, 40, 62, 13, 93, 75] explores producing temporally consistent and controllable video outputs under various manipulation and conditioning tasks. Techniques such as GCD [57], Recapture [87], GS-DiT [8], DAS [19], and recent 4D-consistent pipelines [51, 83, 72] exploit geometric and dynamic scene information, such as tracked 3D points [28, 71], to condition or align the generative models, either via simulation or real-world sequences. These approaches enable synchronizing generated outputs across multiple cameras or time but are typically constrained by how accurately dynamic scene content can be retrieved or tracked. Some works tackle 4D and multi-view video generation by training generative models directly on synchronized video collections [73, 61, 60, 5, 69, 31], or by reconstructing an explicit scene representation first and then rendering views [88, 36, 38, 54].
@@
Ensuring frame-to-frame and cross-view temporal or geometric consistency remains a critical challenge in video synthesis. Early efforts used 3D point clouds or height maps derived from input or generated content to guide learning and enforce consistency [43, 15]. Others [30] propagate consistency across parallel generated video streams but may lose coherence when scene elements leave the views of all sequences. Latent feature histories have also been used to improve consistency for streaming or autoregressive video generation approaches [22], though explicit, interpretable 3D control remains an open research direction.
@@ 3.1 Preliminary: Video Diffusion Models
Our framework builds on the popular architecture used in state-of-the-art text-to-video diffusion models [80, 1, 17, 59], which combine a 3D Variational Auto-Encoder (VAE) with a Diffusion Transformer (DiT) [49]. The VAE spatially and temporally compresses input videos into low-resolution latents that serve as compact representations for diffusion modeling. The diffusion process follows the rectified flow formulation [41], where the model learns to transform noise into coherent video latents through velocity prediction.
@@ 3.2 Network architecture
A naive strategy for injecting camera tokens $\mathbf{c}_{xyt}$ is to simply concatenate or add them to the corresponding video tokens $\mathbf{z}_{xyt}$. A similar approach is adopted in [4], where a camera encoder produces a 12-dimensional pose embedding that is added to the video tokens in every DiT block. However, this formulation entangles the spatial location of the camera and the temporal position of the corresponding frame within the video, as discussed below.
Disentangle camera and temporal position embeddings. Video DiTs use 3D Rotary Positional Embeddings (RoPE) [53] to encode the spatial-temporal positions $(x,y,t)$ of video tokens. Specifically, RoPE applies a frequency-based rotation on the keys and values of a token:
@@ 3.3 Training setup
To enable the model to accommodate different input configurations across a variety of tasks, we leverage a diverse collection of existing 3D/4D datasets encompassing multiple data types. Stereo4D [27] comprises stereo videos with camera-pose annotations; however, because per-video stereo baselines are unavailable, we use a single view with its corresponding pose, effectively treating it as a monocular video dataset. We include RE10K [94] and DL3DV [37] as multi-view image NVS datasets. For image/video NVS, we additionally use the synthetic Syncammaster [5] and Recammaster [4] datasets, featuring multiple time-synchronized static cameras and dynamic camera trajectories, respectively. An overview of the datasets and the tasks each dataset supports is provided in Tab. 1. We primarily target context and target configurations with $1$, 33, 55, or $10$ latent frames, and up to $3$ views for the context input. Owing to the variation in input views and frames across training iterations, we demonstrate in Fig. 5 that our model generalizes to longer video sequences and a greater number of views. Moreover, despite not training directly on the multi-view video NVS task (e.g., $3{\times}3\rightarrow 1{\times}3$), the model generalizes to it naturally, as it can be viewed as a composition of the multi-view image NVS and monocular video NVS tasks.
@@ 4.1 Experimental Setup
Implementation Details. As discussed in Sec. 3.3, we train OmniView on a mixture of the datasets provided in Tab. 1. We choose the Wan 2.1 1.1B model [59] as the base architecture for all our experiments unless mentioned otherwise. Each iteration takes in a set of source views, target views and their corresponding cameras as well as their timestamps in the real world. We train the model for 40K iterations over 32 H100 GPUs. We linearly warmup the learning rate for 3K iterations upto $0.0001$ with a batch size of 64. During the warmup stage, we exclusively train on the multiview static task to quickly update the parameters corresponding to the Plucker ray maps while also allowing the model to quickly adapt to camera conditioning.
@@
TrajectoryCrafter [84] 24.77 3.59 21.36 3.60 14.37 4.00 24.92 2.64 23.11 2.05 21.70 3.18
ReCamMaster [4] 10.18 2.60 11.47 2.19 8.85 3.39 16.49 2.82 13.49 2.37 12.09 2.67
Gen3C [51] 28.38 1.88 42.55 2.04 129.17 2.62 21.80 1.39 25.72 1.71 49.52 1.93
@@
Evaluation and Baselines. We comprehensively evaluate OmniView on a wide variety of 3D/4D tasks. We broadly group multiple tasks under: a) Monocular Video NVS (Tab. 2) b) Multi-view Image and Video NVS (Tab. 3) and c) T2V/I2V + Camera Control. We mainly perform comparisons with closely related generative view-synthesis works of RecamMaster [4], TrajectoryCrafter [83], and GEN3C [51] on the Monocular Video NVS as well as the I2V + Camera Control task. For Multi-view Image NVS, GEN3C and SEVA [91] are our primary baselines which can take multiple view inputs. We visualize the results from our model in Fig. 3 and comparisons in the supplementary. We use the reconstruction metrics of PSNR, SSIM, LPIPS for datasets with GT views available. We additionally evaluate the quality of the camera trajectory via Rotation (RotErr) and Translation Error (TrErr) as proposed by [21]. We use MegaSaM [36] for estimating camera trajectories from generated videos when comparing with the GT trajectories. We additionally compute the metric scale for each scene similar to [2] and scale the camera translation vectors accordingly. We find this necessary for monocular view cases which have inherent scale ambiguity.
@@ 4.2 Monocular Video NVS
For this task, we extract camera trajectories for 45 real-world videos from DAVIS [25]. For each video, we use 4 predefined trajectory paths from RecamMaster evaluations and 1 new trajectory which is a spiral path. Results are summarized in Tab. 2. Averaged over all trajectories, we match the performance of SOTA approaches of [4, 83, 51] in terms of camera error metrics despite being trained on a large diversity of tasks. We additionally visualize qualitative results in the supplementary. We see that we obtain high fidelity generations compared to other works with fewer motion artifacts. We maintain high temporal synchronization compared to the input condition highlighting the capability of the model to accurately extract appearance information from conditioning views while disentangling camera and time conditions.
Beyond camera trajectory error metrics, we also compare with [83, 51] on the Neural 3D Video (N3DV) dataset [33] which consists of multi-view static cameras capturing a dynamic scene. We choose the first view as the test and choose one training view with significant disparity from the test. Compared to prior works, we obtain high quality reconstructions which are well aligned with the GT achieving better PSNR, SSIM and LPIPS. Note that our conditioning is largely implicit via encoded latents and we do not use any form of explicit 3D supervision such as depth or point clouds apart from the metric scene scale. Despite being trained on 122K camera trajectories, RecamMaster fails for this setting of static target camera and tends to implicitly induce camera motion. This highlights the generalizability issue of using 3D spatiotemporal RoPE on camera conditions as discussed in Sec. 3 where the model fails to disentangle camera and time resulting in poor out-of-distribution trajectory performance.